# Direct-to-Cell: A First Look into Starlink’s Direct Satellite-to-Device Radio Access Network through Crowdsourced Measurements

### v6

@@ V Conclusions
The measurements indicate slight increases in RSRP, RSRQ, and SINR throughout the beta testing period, which may be attributable to network changes, including the higher OOBE levels authorized by the FCC and the deployment of some satellites at VLEO altitudes. Based on the SINR measurements, we estimate current average per-beam throughput to be approximately $3.1$ Mbps, which is sufficient for basic services but significantly below terrestrial averages. Looking ahead, per-beam throughput could increase up to $4.2$ Mbps by using VLEO satellites (around $350$ km), and up to $4.5$ Mbps for LEO, assuming that OOBEs relaxation has not been applied yet. In the short term, recent spectrum acquisitions will enable doubling per-beam capacity (to $6.2$ Mbps), with the potential to reach $18.6$ Mbps in the medium term once NTN bands are integrated into commercial smartphone modems. However, the extent to which these enhancements translate into higher per-user throughput will ultimately depend on the expansion of the satellite constellation and the evolution of service demand.

### v7

@@ V Conclusions
The measurements indicate slight increases in RSRP, RSRQ, and SINR throughout the beta testing period, which may be attributable to network changes, including the higher OOBE levels authorized by the FCC and the deployment of some satellites at VLEO altitudes. Based on the SINR measurements, we estimate the current average per-beam throughput to be approximately $3.1$ Mbps, which is sufficient for basic services but significantly below terrestrial averages. Looking ahead, per-beam throughput could increase to $4.2$ Mbps by using VLEO satellites ($\sim 350$ km), and up to $4.5$ Mbps for LEO, assuming that OOBEs relaxation has not been applied yet. In the short term, recent spectrum acquisitions will enable doubling per-beam capacity (to $6.2$ Mbps), with the potential to reach $18.6$ Mbps in the medium term once NTN bands are integrated into commercial smartphone modems. However, the extent to which these enhancements translate into higher per-user throughput will ultimately depend on the expansion of the satellite constellation and the evolution of service demand.